OpenStudy (anonymous):

find the indeical probl if p(a or b) =0.8 p(a)=0.5 and p(a and b)=0.15 find p(b)

OpenStudy (anonymous):

0.45

OpenStudy (watchmath):

P(a or b) = P(a) +P(b)-P(a and b) 0.8 = 0.5+P(b)-0.15 0.8=0.35+P(b) P(b)=0.45.